# A generalisable data-augmented turbulence model with progressive and interpretable corrections

### v1

@@ Abstract
The integration of interpretability and generalisability in data-driven turbulence modelling remains a fundamental challenge for computational fluid dynamics applications. This study yields a generalisable advancement of the Shear Stress Transport (SST) model through a progressive data-augmented framework, combining Bayesian optimisation with physics-guided corrections to improve the predictions of anisotropy-induced secondary flows and flow separation simultaneously. Two interpretable modifications are systematically embedded: 1) a non-linear Reynolds stress anisotropy correction to enhance secondary flow predictions, and 2) an activation-based separation correction in the -equation, regulated by an optimised power-law function to locally adjust turbulent viscosity under adverse pressure gradients. The model is trained using a multi-case computational fluid dynamics-driven a posteriori approach, incorporating periodic hills, duct flow, and channel flow to balance correction efficacy with baseline consistency. Validation across multiple unseen cases – spanning flat-plate boundary layers, high-Reynolds-number periodic hills, and flow over diverse obstacle configurations – demonstrates enhanced accuracy in velocity profiles, recirculation zones, streamwise vorticity, and skin friction distributions while retaining the robustness of the original - SST in attached flows. Sparsity-enforced regression ensures reduced parametric complexity, preserving computational efficiency and physical transparency. Results underscore the framework’s ability to generalise across geometries and Reynolds numbers without destabilising corrections, offering a validated framework toward deployable, data-augmented turbulence models for numerical simulations.
@@ 1 Introduction
The development of turbulence models that can accurately predict complex fluid behaviour while maintaining model consistency, interpretability, and generalisability remains a fundamental challenge in fluid mechanics [1, 2]. Among the Reynolds-Averaged Navier-Stokes (RANS) turbulence models, linear eddy-viscosity models, such as the Spalart-Allmaras [3] and the Shear Stress Transport (SST) models [4], are widely used due to their robustness, computational efficiency, and interpretability. However, these models often fail to capture certain turbulent flow intricacies, particularly in complex geometries and highly unsteady flows [5].
@@
To address these limitations, machine learning techniques have increasingly been integrated into RANS turbulence modelling, offering the potential to enhance accuracy and efficiency while maintaining physical constraints [11, 12]. However, this integration presents significant challenges, particularly in ensuring interpretability and generalisability [13]. In fluid mechanics, model simplicity is crucial, as simpler models tend to favour interpretability and generalisability across diverse flow conditions. Simultaneously, developing a new model that maintains self-similar predictions across flow cases is a complex problem, requiring a balance between predicting universal turbulence features and avoiding overfitting specific cases. Additionally, the inherent limitations of RANS equations in accurately predicting complex turbulence flow physics further complicate the development of stable and generalisable data-driven models. Hence, despite numerous attempts, a universally improved and accepted RANS data-driven turbulence model has yet to be achieved by the scientific community.
@@
Regarding a posterior training, Bin et al. [48, 49, 50] proposed a progressive augmentation approach, which systematically incorporates corrections without compromising baseline model performance. Recent studies by Raje et al. [51] and Oulghelou et al. [52] have likewise prioritised the implementation of blending, flow clustering, and diverse data-driven techniques to reach similar results in two-dimensional cases, although no general, interpretable model nor methodology consensus is achieved to democratise the general use of the findings.
@@
To address the challenges of model consistency, interpretability and generalisability in turbulence modelling, this study employs a progressive augmentation approach to enhance the standard - SST model with data-driven correction factors, specifically targeting anisotropy-induced secondary flows and flow separation phenomena (Fig. 1). The methodology, detailed in Section 2, introduces a correction framework comprising a non-linear Reynolds stress anisotropy modification embedded within the velocity transport equation and an activation-based correction in the -equation to refine turbulence behaviour under adverse pressure gradients. Model training is conducted using an a posteriori multi-case CFD-driven optimisation framework, incorporating duct flow, periodic hills, and channel flow cases to ensure a balanced correction strategy. Section 3 presents the verification and validation of the trained model on unseen test cases, encompassing a range of Reynolds numbers and geometric configurations. Finally, Section 4 summarises the findings and discusses the implications of the proposed framework for data-driven turbulence modelling. By systematically ensuring model consistency, interpretability, and generalisability, this work advances the integration of physics-informed data augmentation within RANS turbulence modelling, ensuring reliability across a broad range of flow conditions while preserving physical interpretability.
@@ 2.1 Reynolds stress anisotropy correction
To improve secondary flow predictions, the linear eddy-viscosity model (Eq. 4) is extended using Pope’s decomposition [60] and adding the non-linear normalised second tensor with an unknown correction function as follows:
where is the mean rotation-rate tensor. The addition of these two variables ( and ), ensures the preservation of the canonical model’s turbulent viscosity prediction while only enhancing Reynolds stress anisotropy, and therefore, Prandtl second type of secondary flow.
@@ 2.2 Flow separation correction
This approach does not come from a new design of machine learning but is derived from a detailed analysis of the governing equations, ensuring that the correction remains consistent with the underlying physical principles. The correction term directly targets regions where conventional linear eddy-viscosity turbulence models struggle, particularly under adverse pressure gradient conditions where deviations from equilibrium assumptions occur [59]. To ensure that the correction is both effective and selective, a power-law activation function, , is employed, following
where and are additional optimisation parameters to ensure the accurate implementation of the separation correction (Fig. 2). This function regulates the activation of the correction term, ensuring it only engages where local corrections are necessary, thus preserving the model’s predictive consistency in attached flows.
@@ 2.5 Training and validation cases
To date, there has not been a standardised method for thoroughly validating data-driven turbulence models. However, various studies have utilised a set of CFD cases, such as flat-plate boundary layer development and flow over different obstacles, to assess model performance. In this study, the generalisation capabilities of the trained model are tested across different geometries, Reynolds numbers, and flow characteristics based on prior literature. Hence, the training cases consist of a duct flow case at bulk Reynolds number of 3500 and aspect ratio of 1 (DF), flow over periodic hills at bulk Reynolds number of 3500 (PH2800), and channel flow at frictional Reynolds number of 590 (CF590). Testing cases span from diverse Reynold numbers from the training cases to different geometries as depicted in Table 1. Notably, the model is tested against the flat-plate FP, and wall-mounted hump WMH [73, 74, 75] cases, included in the 2022 NASA Symposium on Turbulence Modelling [76]. This initiative challenged the research community to develop and evaluate state-of-the-art data-driven turbulence models, addressing the needs of the field while identifying potential pitfalls and strategies to improve model generalisability and performance.
@@ 3.1 Training cases results
Qualitatively, Fig. 6(a) shows that the PDA-enhanced model improves the prediction of streamwise velocity profiles, especially in the vicinity of the duct walls where secondary flow effects are most pronounced. The baseline model completely neglects any secondary motion, whereas the PDA model corrects this by predicting RST anisotropy more accurately. Quantitatively, in Fig. 6(b), the PDA model demonstrates a notable enhancement in predicting the characteristic in-plane vortical structures. The baseline model does not predict a secondary flow field, whereas the PDA correction is capable of doing so, amplifying the in-plane flow effects and more accurately predicting the high-fidelity secondary flow. Importantly, although the separation correction is active, the corrections remain inactive for this case adequately, ensuring that improvements are selectively applied to enhance anisotropy representation without introducing spurious modifications elsewhere.
@@ 3.2 Validation data
To further assess the generalisability and the separation correction , five validation cases are considered: a higher Reynolds number case PH10595 and four additional cases – CBFS13700, BUMP20, BUMP42, and NASA challenge case WMH. For conciseness, only the detailed results for CBFS13700 and WMH are presented here, while final results for PH10595, BUMP20, and BUMP42 are provided in Table 3.
@@
The friction and pressure coefficient distributions for case WMH are presented in Fig. 11. Despite this case not being included in the training set, a consistent improvement is observed, aligning with the trends seen in other validation cases. A slight overprediction of is observed at (Fig. 11(a)), where the experimental data exhibit a higher degree of uncertainty – an effect also noted in the BUMP cases. Nevertheless, the recovery of following separation closely matches the experimental data, showing a significant improvement. For the pressure coefficient () (Fig. 11(b)), the results remain largely consistent with the predictions of the standard - SST model. However, slight improvements are observed at , where the PDA-enhanced predictions align more closely with the experimental data, further reinforcing the model’s ability to enhance flow recovery dynamics. Additionally, friction coefficient distributions reveal that the baseline model underestimates the magnitude of in the separation region and fails to capture the wall-shear stress recovery downstream accurately. The PDA model corrects this behaviour by locally increasing , reducing the overprediction of the recirculation length, and improving shear stress gradients for all cases.
@@
To establish a final validation assessment, the model is further examined where the anisotropy and separation corrections are not expected to influence the flow predictions. In these cases, the specific physical effects that the model’s corrections are designed to address are absent. Consequently, the PDA-enhanced formulation should yield results that are identical to the standard - SST model. Any significant deviation would indicate unintended modifications to well-predicted flow regions. In order to assess this influence, the FP case is evaluated. As shown in Fig. 12, the PDA model produces predictions that are indistinguishable from those of the standard - SST model for both the streamwise velocity and the friction coefficient. This outcome further demonstrates the robustness of the activation function , which effectively prevents corrections from being applied in regions where they are unnecessary, ensuring that the baseline model’s accuracy is preserved.
| -1.584 | -0.685 | -0.178 | -2.070 | 1.119 | -0.215 | 18.622 | 4.698 |
These results validate the PDA framework’s ability to enhance separation and anisotropy-based secondary flow predictions while preserving the robustness of the original model, addressing a critical limitation of the - SST formulation for adverse pressure gradient flows and non-linear Reynolds stress corrections. Finally, Table 4 shows the final coefficient values optimised for the model.

### v2

@@ Abstract
The integration of interpretability and generalisability in data-driven turbulence modelling remains a fundamental challenge for computational fluid dynamics applications. This study yields a generalisable advancement of the - Shear Stress Transport (SST) model through a progressive data-augmented framework, combining Bayesian optimisation with physics-guided corrections to improve the predictions of anisotropy-induced secondary flows and flow separation simultaneously. Two interpretable modifications are systematically embedded: 1) a non-linear Reynolds stress anisotropy correction to enhance secondary flow predictions, and 2) an activation-based separation correction in the -equation, regulated by an optimised power-law function to locally adjust turbulent viscosity under adverse pressure gradients. The model is trained using a multi-case computational fluid dynamics-driven a posteriori approach, incorporating periodic hills, duct flow, and channel flow to balance correction efficacy with baseline consistency. Validation across multiple unseen cases – spanning flat-plate boundary layers, high-Reynolds-number periodic hills, and flow over diverse obstacle configurations – demonstrates enhanced accuracy in velocity profiles, recirculation zones, streamwise vorticity, and skin friction distributions while retaining the robustness of the original - SST in attached flows. Sparsity-enforced regression ensures reduced parametric complexity, preserving computational efficiency and physical transparency. Results underscore the framework’s ability to generalise across geometries and Reynolds numbers without destabilising corrections, offering a validated framework toward deployable, data-augmented turbulence models for numerical simulations.
@@ 1 Introduction
The development of turbulence models that can accurately predict complex fluid behaviour while maintaining model consistency, interpretability, and generalisability remains a fundamental challenge in fluid mechanics [1, 2]. Among the Reynolds-Averaged Navier-Stokes (RANS) turbulence models, linear eddy-viscosity models, such as the Spalart-Allmaras [3] and the Shear Stress Transport (SST) models [4], are widely used due to their robustness, computational efficiency, and interpretability. However, these models often fail to capture certain turbulent flow intricacies, particularly in complex geometries and highly unsteady flows [5, 6].
@@
To address these limitations, machine learning techniques have increasingly been integrated into RANS turbulence modelling, offering the potential to enhance accuracy and efficiency while maintaining physical constraints [12, 13]. However, this integration presents significant challenges, particularly in ensuring interpretability and generalisability [14]. The turbulence modelling community widely recognises that achieving a fully generalisable data-driven model (one that performs reliably across all flow physics and geometries) remains an extraordinary undertaking and a long-term goal [15, 16, 2, 17, 18]. Progress towards this objective often involves methodological advancements that demonstrate robust performance improvements for specific, challenging classes of flows. A key challenge is therefore to develop frameworks that allow for targeted, non-detrimental augmentation for distinct physical phenomena, validated on a range of unseen but related cases. This controlled demonstration of generalisability for targeted improvements, without compromising the baseline model’s performance where it is already reliable, is a critical step forward and one that many current data-driven approaches find challenging.
@@
Regarding a posterior training, Bin et al. [53, 54, 55] proposed a progressive augmentation approach, which systematically incorporates corrections without compromising baseline model performance. Recent studies by Raje et al. [56] and Oulghelou et al. [57] have likewise prioritised the implementation of blending, flow clustering, and diverse data-driven techniques to reach similar results in two-dimensional cases, although no general, interpretable model or methodology consensus is achieved to democratise the general use of the findings.
@@
Hence, to address the challenges of model consistency, interpretability and generalisability in turbulence modelling, this study employs a progressive augmentation approach to enhance the standard - SST model with data-driven correction factors, specifically targeting anisotropy-induced secondary flows and flow separation phenomena (Fig. 1). The methodology, detailed in Section 2, introduces a correction framework comprising a non-linear Reynolds stress anisotropy modification embedded within the velocity transport equation and an activation-based correction in the -equation to refine turbulence behaviour under adverse pressure gradients. Model training is conducted using an a posteriori multi-case CFD-driven optimisation framework, incorporating duct flow, periodic hills, and channel flow cases to ensure a balanced correction strategy. Section 3 presents the verification and validation of the trained model on unseen test cases, encompassing a range of Reynolds numbers and geometric configurations. Finally, Section 4 summarises the findings and discusses the implications of the proposed framework for data-driven turbulence modelling. The proposed PDA framework is developed following established guidelines for data-driven turbulence modelling, such as those defined by Spalart [15, 16]. The new model aims to leverage the strengths of a well-validated baseline RANS model ( SST) by introducing physics-informed, interpretable corrections with a limited number of optimisable parameters. The a posteriori training methodology ensures numerical stability and solver-consistency, while the design of the correction terms, particularly an activation function for the separation correction, aims to preserve the baseline model’s accuracy in well-predicted flow regimes and ensure robust behaviour. This approach consciously tries to avoid pitfalls associated with purely data-driven black-box models and prioritises generalisability and physical consistency. By systematically ensuring model consistency, interpretability, and generalisability, this work advances the integration of physics-informed data augmentation within RANS turbulence modelling, ensuring reliability across a broad range of flow conditions while preserving physical interpretability.
It is important to clarify the scope of the present investigation. The developed framework and the resulting model are specifically focused on improving predictions for incompressible, internal and wall-bounded flows. The data-driven corrections are tailored to address deficiencies observed within this flow class. For instance, the under-estimation of turbulent viscosity in the separated internal flow cases studied here is a known issue that the proposed correction targets. This behaviour is notably distinct from that observed in many external aerodynamic applications, where standard RANS models tend to over-predict turbulent viscosity, leading to delayed stall predictions. Consequently, the direct application of the present model is confined to the class of flows for which it was developed. The extension of this progressive augmentation methodology to other regimes, such as external aerodynamics or compressible flows, represents a valuable direction for future work but is beyond the scope of this study.
@@ 2.1 Reynolds stress anisotropy correction
To improve secondary flow predictions, the linear eddy-viscosity model (Eq. 4) is extended using Pope’s decomposition [65] and adding the non-linear normalised second basis tensor with an unknown correction function as follows:
where is the mean rotation-rate tensor. Specifically, this modified is incorporated into the mean momentum transport equations (Eq. 4). Subsequently, the turbulent production term, , is calculated using this modified . Finally, this production term , now influenced by the correction, is then used as the production term in both the -transport equation and the -transport equation. The addition of these two variables ( and ) ensures the preservation of the canonical model’s turbulent viscosity prediction while only enhancing Reynolds stress anisotropy, and therefore, Prandtl’s second type of secondary flow. While a full tensor basis expansion for 2D flows includes an additional term to fully represent the Reynolds-stress anisotropy [65, 66], this study focuses only on the integration of the second tensor basis due to its established role in predicting anisotropy-induced secondary flows [58, 59], and to manage computational optimisation costs. Future work could explore the inclusion of further bases.
@@ 2.2 Flow separation correction
This approach is not a conventional machine-learning design but is derived from a detailed analysis of the governing equations, ensuring consistency with underlying physical principles. The correction term directly targets regions where linear eddy-viscosity models struggle, particularly under adverse pressure gradients where equilibrium assumptions fail [64]. To ensure the correction is both effective and selective, the activation function is designed to act as a switch, engaging the correction term only where physically necessary. The design is based on the behaviour of the baseline SST model itself, particularly the relationship between turbulent viscosity , turbulent kinetic energy , and specific dissipation rate . As observed by Menter [4], in well-behaved equilibrium boundary layers, . Under these conditions, the correction should be inactive. Conversely, in regions of strong adverse pressure gradients and flow separation, the SST model’s limiters often cause to be significantly smaller than . To leverage this behaviour, a power-law activation function is defined as follows:
where and are additional optimisation parameters. This function is monotonically decreasing with respect to the term . When the flow is attached, , which drives towards zero, effectively deactivating the correction and recovering the standard model’s behaviour. However, in separated regions where , becomes positive, switching on the correction. The exponents and are optimised to control the sharpness and strength of this transition, ensuring a numerically stable activation. This physics-guided design preserves the model’s predictive consistency in attached flows while targeting its known deficiencies in separated flows. This behaviour can be seen in more detail in Fig. 2.
@@ 2.5 Training and validation cases
To date, there has not been a standardised method for thoroughly validating data-driven turbulence models. However, various studies have utilised a set of CFD cases, such as flat-plate boundary layer development and flow over different obstacles, to assess model performance. In this study, the generalisation capabilities of the trained model are tested across different geometries, Reynolds numbers, and flow characteristics based on prior literature. Hence, the training cases consist of a duct flow case at bulk Reynolds number of 3500 and aspect ratio of 1 (DF), flow over periodic hills at bulk Reynolds number of 2800 (PH2800), and channel flow at frictional Reynolds number of 590 (CF590). Testing cases span from diverse Reynold numbers from the training cases to different geometries as depicted in Table 1. Notably, the model is tested against the flat-plate FP, and wall-mounted hump WMH [78, 79, 80] cases, included in the 2022 NASA Symposium on Turbulence Modelling [81]. This initiative challenged the research community to develop and evaluate state-of-the-art data-driven turbulence models, addressing the needs of the field while identifying potential pitfalls and strategies to improve model generalisability and performance.
@@ 2.7 Training
A key aspect of this a posteriori training methodology is its inherent robustness against overfitting. In this study, overfitting has not been observed in the optimisation runs, which is attributed to the requirement of achieving a converged, steady-state solution for multiple, physically distinct flow scenarios simultaneously. Nonetheless, the need to achieve a converged steady-state solution inherently limits the improvement thresholds that can be achieved, particularly for the non-linear corrections, whose non-linear nature can destabilise the numerical model. Moreover, the inclusion of the channel-flow case (CF590) in all optimisation objectives further mitigates overfitting, since any aggressive optimisation towards high-fidelity data in the other cases that would negatively impact the law-of-the-wall prediction is inherently penalised. Hence, each training observation in our methodology incorporates three distinct cases (DF3500,AR=1, PH2800, and CF590), a multi-case strategy which discourages overfitting and promotes a balanced, generalisable model.
@@ 3.1 Training cases results
Qualitatively, Fig. 6(a) shows that the PDA-enhanced model improves the prediction of streamwise velocity profiles, especially in the vicinity of the duct walls where secondary flow effects are most pronounced. The baseline model completely neglects any secondary motion, whereas the PDA model corrects this by predicting RST anisotropy more accurately. Quantitatively, in Fig. 6(b), the PDA model demonstrates a notable enhancement in predicting the characteristic in-plane vortical structures. The baseline model does not predict a secondary flow field, whereas the PDA correction is capable of doing so, amplifying the in-plane flow effects and more accurately predicting the high-fidelity secondary flow. Importantly, although the separation correction is active, the corrections remain inactive for this case, adequately ensuring that improvements are selectively applied to enhance anisotropy representation without introducing spurious modifications elsewhere. It is important to highlight the slight discrepancies between the improved model and the high-fidelity data. Since this is a training case, one could expect that the model’s predictions would match the training data, however, due to the pursue of a stable steady-state solution in the CFD solver and the simplification of the full tensor basis expansion for 2D flows (including only ), the final improvement achieved in the a posteriori training performed is constrained to match the high-fidelity data fully.
@@
| -1.584 | -0.685 | -0.178 | -2.070 | 1.119 | -0.215 | 18.622 | 4.698 |
It is noted that a discrepancy in skin friction persists in the attached flow region on the crest of the hill, where both the baseline and the PDA model over-predict the high-fidelity data. This behaviour is a known characteristic of RANS models applied to this periodic flow case. The PDA model is designed as a targeted correction for separated flows, and its activation function, , is explicitly intended to confine corrections to regions where the baseline model struggles most. The observation that the PDA model’s prediction mirrors the baseline SST in this attached region is therefore not a failure, but rather a confirmation that the activation function is operating as intended. Furthermore, the mesh resolution used for the a posteriori RANS optimisation is necessarily far coarser than that of the reference DNS [84]; achieving local DNS-level accuracy on a RANS grid is computationally prohibitive and not the objective of this framework. The primary goal (improving the prediction of the downstream separation) is successfully achieved, demonstrating a robust and practical advancement.
To better illustrate the spatial impact and interpretable nature of the PDA model’s corrections, Fig. 8 displays contours of the key correction components for both the DF and PH cases and Table 4 shows the final coefficient values optimised for the model. For the duct flow, it is observed that the flow invariants and exhibit distinct anticorrelated spatial patterns in the corner regions. When combined to form the anisotropy correction , the resulting spatial structure strongly resembles that of the second invariant, , highlighting its importance in capturing the features of the secondary flow. For the periodic hill case, which drives the separation correction, the invariants are concentrated in the shear layer detaching from the hill crest. Here, the structure of the resulting separation correction, , is clearly dominated by the contribution from , which is consistent with the larger magnitude of its corresponding coefficient, , in Table 4. This analysis visually demonstrates how the framework uses physically-meaningful, spatially-varying inputs to build targeted corrections for specific flow phenomena.
The mathematical structure of the corrections and , as linear combinations of invariants plus a constant term, is explicit and transparent. The inputs to these corrections, specifically the invariants , are derived from well-defined, physically meaningful local flow quantities (); while the activation function is based on quantities () that characterise the state of turbulence and its deviation from equilibrium. The coefficients (), while determined through data-driven optimisation, act as fixed weights for these physically interpretable terms. The final values of the coefficients and (presented in Table 4) provide an indication of the learned importance and direction of influence of the respective standardised invariants on the corrections and . This approach is distinct from less transparent models (where the entire correction might be, for example, the direct output of a complex neural network).
@@ 3.2 Validation data
To further assess the generalisability and the separation correction , five validation cases are considered: a higher Reynolds number case PH10595 and four additional cases — CBFS13700, BUMP20, BUMP42, and NASA challenge case WMH. For conciseness, only the detailed results for CBFS13700 and WMH are presented here, while final results for PH10595, BUMP20, and BUMP42 are provided in Table 3 and A.
@@
To ensure a direct and fair comparison, the simulations were initialised by interpolating the velocity profiles from the high-fidelity reference data directly at the inlet plane. This is why the skin-friction values for all models match the high-fidelity data precisely at the start of the domain, as seen in Fig. 11(b). However, a RANS model has its own internal equilibrium for a turbulent boundary layer, which naturally differs from that of the DNS. Hence, immediately downstream of the inlet, the SST model adjusts the flow from the imposed state towards its own equilibrium, creating a difference in results for comparwed to high-fidelity data. This method was chosen deliberately to guarantee that all simluations for this case start from the exact same conditions, thereby isolating the performance differences purely to the turbulence models themselves.
The wall-mounted hump case (WMH) is included as a particularly challenging test of the model’s extrapolative capabilities. This case is a well-established benchmark within the turbulence modelling community, partly due to documented difficulties in reproducing the exact experimental conditions, particularly in the upstream attached flow, where simulations often mismatch experimental skin friction [81]. The PDA model’s separation correction is designed primarily to improve the local physics of the separation bubble and reattachment phenomena, most directly reflected in the skin friction () distribution. The pressure coefficient (), being a more globally-influenced quantity, is expected to be less sensitive to such localised corrections.
The results for this case are presented in Fig. 12. Despite the case’s inherent complexities and the noted upstream mismatch, the PDA model shows a consistent improvement where the correction is designed to be active. The baseline model’s underestimation of the magnitude in the separation region and its failure to capture wall-shear stress recovery are significantly improved by the PDA correction. This is achieved by locally increasing , which reduces the overprediction of the recirculation length. The recovery of following separation closely matches the experimental data, demonstrating that the local correction is working as intended. For the pressure coefficient () (Fig. 11(b)), the results remain largely consistent with the predictions of the standard - SST model, though slight improvements are observed at , where the PDA-enhanced predictions align more closely with the experimental data. Evaluating the model against this complex case thus provides a valuable demonstration of the correction’s targeted effectiveness and robustness.
@@
To establish a final validation assessment, the model is further examined where the anisotropy and separation corrections are not expected to influence the flow predictions. In these cases, the specific physical effects that the model’s corrections are designed to address are absent. Consequently, the PDA-enhanced formulation should yield results that are identical to the standard - SST model. Any significant deviation would indicate unintended modifications to well-predicted flow regions. In order to assess this influence, the FP case is evaluated. As shown in Fig. 13, the PDA model produces predictions that are indistinguishable from those of the standard - SST model for both the streamwise velocity and the friction coefficient. To compare the friction coefficient, White’s [86] correlation is used. This correlation has been studied by Nagib et al. (2007) [87] and showed great agreement between experimental and high-fidelity numerical data. This outcome further demonstrates the robustness of the activation function , which effectively prevents corrections from being applied in regions where they are unnecessary, ensuring that the baseline model’s accuracy is preserved.
These results validate the PDA framework’s ability to enhance separation and anisotropy-based secondary flow predictions while preserving the robustness of the original model, addressing a critical limitation of the - SST formulation for adverse pressure gradient flows and non-linear Reynolds stress corrections.